OpenStudy (tina1415):

A painter uses 2/3 gallon of paint on one job and 3/5 of the same paint on another. How many gallons of paint did the painter use for both jobs?

OpenStudy (anonymous):

Its 2/3 + (1 - 2/3)*3/5 = 2/3 + 1/5 = 13/15